OpenStudy (anonymous):

A basketball team consists of five players— one centre, two forwards, and two guards. The senior squad at Vennville Central High School has two centres, six forwards, and four guards. c) How many of these starting lineups include Dana, the team’s 185-cm centre? d) Some coaches designate the forwards as power forward and small forward. If all six forwards are adept in either position, how would this designation affect the n

OpenStudy (kropot72):

c) There are 6C2 combinations of forwards and 4C2 combinations of guards. The total number of combinations of forwards and guards each of which could include Dana is 6C2 * 4C2.\[6C2\times 4C2=\frac{6!}{4!2!}\times \frac{4!}{2!2!}=You\ can\ calculate\]